ghostrider:

A circle has its center at (−2, −3) and a radius of 3 units. What is the equation of the circle?

XioGonz:

Do you have an idea of what your equation might look like?

ghostrider:

(x + 2)2 + (y + 3)2 = 3 here is what I think

XioGonz:



darkknight:

Almost, except \[x-h)^2+(y-k)^2=r^2\] is formula for center of circle

darkknight:

so if r =3. What is r^2?

XioGonz:

K,..takeover.

ghostrider:

r^2 is 9

darkknight:

then your equation is (x + 2)2 + (y + 3)2 = 9 to fit the formula of a circle

darkknight:

You see why not right?

ghostrider:

thanks yea I do

darkknight:

alright np, have a good day